# A Host–Kra $\mathbb{F}_{2}^{\omega}$-system of order $5$ that is not Abramov of order $5$, and non-measurability of the inverse theorem for the $U^{6}(\mathbb{F}_{2}^{n})$ norm

### v2

@@ 1.2. Acknowledgments
OS was supported by NSF grant DMS-1926686 and ISF grant 3056/21. TT was supported by a Simons Investigator grant, the James and Carol Collins Chair, the Mathematical Analysis & Application Research Fund Endowment, and by NSF grant DMS-1764034. We thank Tim Austin, Pablo Candela, and Luka Milićević for corrections and comments.
@@ 2. A characterization of coboundaries on $\mathbb{F}_{2}^{n}$
where $h_{i}=(h_{i}^{(1)},h_{i}^{(2)})\in G$, one can verify that $\rho$ is a $2$-cocycle on the elementary abelian $2$-group $G$ that does not obey (7). This $2$-cocycle will be related to a non-trivial (but now $2$-homogeneous) $5$-cocycle on the degree $2$ filtered abelian group ${\mathcal{D}}^{2}(\mathbb{F}_{2}^{2})$ that we will construct in the next section.
@@
A notable special case of this theorem occurs when $\rho_{h_{1},\dots,h_{k}}$ is constant for each $h_{1},\dots,h_{k}$, then the $2$-homogeneous $k$-cocycle $\rho$ is what is referred to as a non-classical symmetric multilinear form in [27] and a strongly symmetric multilinear form in [21], and the potential $F$ produced by this theorem is then a (non-classical) polynomial of degree $k$. This special case of Theorem 2.3 was previously established in [27, Proposition 3.5].
@@
Now suppose inductively that $k>0$ and the claim has already been proven for $k-1$. For each $h_{1}\in G$, the function $\rho_{h_{1}}\colon G^{[k]}\to\mathbb{T}$ defined by $(\rho_{h_{1}})_{h_{2},\dots,h_{k+1}}(x)\coloneqq\rho_{h_{1},\dots,h_{k+1}}(x)$ can be easily verified to be a $2$-homogeneous $(k-1)$-cocycle. Hence by the induction hypothesis, there exists $F_{h_{1}}\colon G\to\mathbb{T}$ such that
@@
It remains to establish (12). We prove this by a further induction on the dimension $n$. The case $n=0$ is trivial, so suppose $n\geq 1$ and that the claim has already been proven for $n-1$. Now split $G=\mathbb{F}_{2}^{n-1}\times\mathbb{F}_{2}$ and let $e=(0,1)$ be the generator for the $\mathbb{F}_{2}$ factor. The operator $\partial_{e}$ is annihilated by $1+T^{e}$ since $(1+T^{e})\partial_{e}=\partial_{2e}=0$. Also, for $k>2$, the operator $1+T^{e}=2+\partial_{e}$ maps $\mathrm{Poly}^{k-2}(G)$ to $\mathrm{Poly}^{k-3}(G)$ thanks to (61), hence from (10)
@@
Since $F_{e}-F^{\prime}_{e}$ is annihilated by $1+T_{e}$, we may write
@@
for some $\phi\colon G\to\mathbb{T}$. If we then write
@@
is a polynomial of degree $k-2$ on $G$, while
@@
for all $h\in\mathbb{F}_{2}^{n-1}\times\{0\}$. On the other hand, from (9), the vanishing of $F^{\prime\prime}_{e}$, and the $e$-invariance of $\phi^{\prime\prime}$ we see that
@@
First suppose that $\rho$ is a $k$-coboundary in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, thus $\rho=d^{k+1}F$ for some $F\colon G\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$. Then we can write $\rho=d^{k}\psi$ with $\psi\coloneqq dF$; since $02F=$ 0, we have $02\psi=$ 0, and so $\rho$ is certainly strongly $2$-homogeneous.
@@ 3. A non-trivial cocycle
There exists a strongly $2$-homogeneous $5$-cocycle $\rho\colon C^{6}(X_{2})\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ on $X_{2}$ taking values in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, which is not a $5$-coboundary (when viewed as a cocycle in $\mathbb{T}$).
@@
It will be convenient to adopt the following notation from [26, Definitions 6.1, 6.3].
@@
$\quad\coloneqq\sum_{\{\{i_{1},\dots,i_{k}\},\{j_{1},\dots,j_{k}\}\}:\{1,\dots,2k\}=\{i_{1},\dots,i_{k}\}\cup\{j_{1},\dots,j_{k}\}}S(h_{i_{1}},\dots,h_{i_{k}})S(h_{j_{1}},\dots,h_{j_{l}}).$
@@
and $B*B=2\mathrm{Sym}^{2}(B)$; in particular, in characteristic two we have $B*B=0$. The trilinear form in Remark 2.2 can be written as
(17) $\rho_{h_{1},h_{2},h_{3}}(x)=(\mathrm{Sym}^{2}(L_{1})*L_{2})(h_{1},h_{2},h_{3})$
@@
A $6$-cube in $X_{2}={\mathcal{D}}^{2}(\mathbb{F}_{2}^{2})$ can be viewed as a pair $(Q^{(1)},Q^{(2)})$, where $Q^{(1)},Q^{(2)}\colon\mathbb{F}_{2}^{6}\to\mathbb{F}_{2}$ are quadratic polynomials, so in particular their second derivatives can be viewed as symmetric bilinear forms $d^{2}Q^{(i)}\colon\mathbb{F}_{2}^{6}\times\mathbb{F}_{2}^{6}\to\mathbb{F}_{2}$, defined for $i=1,2$ by the formula
@@
with $e_{1},\dots,e_{6}$ the standard basis of $\mathbb{F}_{2}^{6}$; compare with (17).
@@
The function $\rho$ is clearly symmetric with respect to permutations of the indices $1,\dots,6$. If we fix the $h_{ij}$ for $1<i<j\leq 6$, then this function is linear in the remaining variables $h_{1i}$, $61<i<$ 6, from which it is easy to verify that $\rho$ obeys the cocycle property in Definition A.6(ii). Thus $\rho$ is a $5$-cocycle.
@@
Finally, we need to show that $\rho=d^{5}\psi$ for some $\psi\colon X_{2}^{[1]}\to\mathbb{T}$ with $2\psi$ a cubic polynomial. We let $[]\colon\mathbb{F}_{2}\to\mathbb{Z}/4\mathbb{Z}$ be any left inverse of the projection map $\mod 2\colon\mathbb{Z}/4\mathbb{Z}\to\mathbb{F}_{2}$; in particular one has $[0]^{2}=0\mod 4$ and $[1]^{2}=1\mod 4$ regardless of the choice of left inverse. An element of $C^{1}(X_{2})$ takes the form $(x,x+h)$ with $x,h\in\mathbb{F}_{2}^{2}$. We write $x=(x^{(1)},x^{(2)})$, $h=(h^{(1)},h^{(2)})$ and define $\psi$ by the formula
$\psi(x,x+h)\coloneqq\frac{[x^{(1)}]^{2}[h^{(2)}]^{2}}{4}+\frac{x^{(1)}h^{(1)}x^{(2)}}{2}\mod 1.$
We first verify that $2\psi$ is a cubic polynomial. Since $[x]^{2}=x^{2}=x\mod 2$, we have
$2\psi(x,x+h)=\frac{x^{(1)}h^{(2)}}{2}\ \mod 1$
and a $4$-cube in $C^{1}(X_{2})$ takes the form
@@
for some $x,h_{0},h_{i},h_{0i},h_{ij}\in X_{2}$. The function $d^{4}(2\psi)$ applied to this cube is then equal to
@@
It remains to show that $\rho=d^{5}\psi$. A $5$-cube in $X_{2}^{[1]}$ takes the form
@@
for some $x,h_{0},h_{i},h_{0i},h_{ij}\in X_{2}$. The function $d^{5}\psi$ applied to this cube is the sum of
@@
for $a=1,2$. We first consider (23). The numerator $X^{(1)}(\omega)H^{(1)}(\omega)X^{(2)}(\omega)$ is quintic in the $\omega_{i}$, so the alternating sum $\sum_{\omega\in\{0,1\}^{5}}(-1)^{5-|\omega|}$ is extracting the $\omega_{1}\dots\omega_{5}$ coefficient of this numerator, which after expanding out all the definitions can be expressed as
@@
where the sum $\sum^{*}$ ranges over the $30$ pairs of sets $\{\{a,b\},\{c,d\}\},\{e,f\}$ with $\{0,1,2,3,4,5\}=\{a,b\}\cup\{c,d\}\cup\{e,f\}$ such that 00 lies in one of $\{a,b\}$ or $\{c,d\}$.
Now consider (22). Using the easily verified identities $[a+b]^{2}=[a]^{2}+[b]^{2}+2[ab]$ and $[a\omega]^{2}=[a]^{2}\omega$ for $a,b\in\mathbb{F}_{2}$ and $\omega\in\{0,1\}$ (and noting that the map $a\mapsto 2[a]$ is an additive homomorphism), we can expand out
$[H^{(2)}(\omega)]^{2}=\left[h^{(2)}_{0}+\sum_{i=1}^{5}h^{(2)}_{0i}\omega_{i}\right]^{2}=[h^{(2)}_{0}]^{2}+\sum_{i=1}^{5}[h^{(2)}_{0i}]^{2}\omega_{i}+2[Q(\omega)]$
@@
$[X^{(1)}(\omega)]^{2}=[x^{(1)}]^{2}+\sum_{i=1}^{5}[h^{(1)}_{i}]^{2}\omega_{i}+\sum_{1\leq i<j\leq 5}[h^{(1)}_{ij}]^{2}\omega_{i}\omega_{j}+2[R(\omega)]$
@@
The product $[X^{(1)}(\omega)]^{2}[H^{(2)}(\omega)]^{2}$ is then quintic (the product of $2Q$ and $2R$ would be sextic, but vanishes modulo $4$), and the alternating sum $\sum_{\omega\in\{0,1\}^{5}}(-1)^{5-|\omega|}$ is then extracting the $\omega_{1}\dots\omega_{5}$ coefficient, which can only arise from the terms
@@
in the numerator. Inspecting the cubic terms of $R(\omega)$, we conclude that (23) is of the form
@@
where the sum $\sum^{**}$ ranges over the $15$ pairs of sets $\{\{a,b\},\{c,d\}\},\{e,f\}$ with $\{0,1,2,3,4,5\}=\{a,b\}\cup\{c,d\}\cup\{e,f\}$ such that 00 does not lie in either $\{a,b\}$ or $\{c,d\}$. Summing (24), (25), we obtain the claim. This concludes the proof of Theorem 3.1.
@@ 4. Two key nilspaces
We now use the non-trivial cocycle $\rho$ introduced in the previous section to construct family of finite $5$-step nilspaces $X_{5,r}$ for $1\leq r\leq 5$ that will play a key role in our counterexamples. To prove our main result in Theorem 1.6 we will use the larger and more complicated nilspace $X_{5,5}$, however in Proposition 4.5 below we obtain a weaker counterexample with significantly less effort using the smaller and simpler nilspace $X_{5,1}$.
@@
We first show existence. Since $\rho$ is a strongly $2$-homogeneous $5$-cocycle, it is not difficult to see that the pullback $Q^{*}\rho$ is also. Hence by Theorem 2.6, $Q^{*}\rho$ is a $5$-coboundary in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, thus there exists $S\colon\mathbb{F}_{2}^{n}\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ such that $Q^{*}\rho=d^{6}S$, which is precisely the condition (1). Since $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ is contained in $\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$, we have obtained an $n$-cube $(Q,S)$ in $X_{5,r}$ as required.
@@
In fact, with the specific choice of cocycle we have constructed, we can describe the coset in Lemma 4.1 explicitly.
@@
Let the notation and hypotheses be as in Lemma 4.1. Write $Q=(Q_{1},Q_{2})$, thus $Q_{1},Q_{2}\colon\mathbb{F}_{2}^{2}\to\mathbb{F}_{2}$ are classical quadratic polynomials. Let $R\in\mathrm{Poly}^{3}(\mathbb{F}_{2}^{2}\to\mathbb{Z}/4\mathbb{Z})$ be a cubic polynomial such that $2\frac{R}{4}=\frac{Q_{1}}{2}\mod 1$ (or equivalently that $R=Q_{1}\mod 2$); the existence of such a polynomial follows from (61). Then the coset of $S$ in Lemma 4.1 is equal to
@@
where (as in Section 1.3) $\binom{a}{2}\in\mathbb{F}_{2}$ is equal to $1$ when $a=2,3\mod 4$ and 00 for $a=0,1\mod 4$.
@@
for $x\in\mathbb{F}_{2}^{n}$ and $\vec{h}=(h_{1},\dots,h_{6})\in(\mathbb{F}_{2}^{n})^{6}$. By construction of $\rho$, it suffices to show that
$\partial_{h_{1}}\dots\partial_{h_{6}}\left(\binom{R}{2}Q_{2}\right)=\sum_{\{\{a,b\},\{c,d\}\},\{e,f\}:\{1,\dots,6\}=\{a,b\}\cup\{c,d\}\cup\{e,f\}}(\partial_{h_{a}}\partial_{h_{b}}Q_{1})(\partial_{h_{c}}\partial_{h_{d}}Q_{1})(\partial_{h_{e}}\partial_{h_{f}}Q_{2})$
for all $h_{1},\dots,h_{6}\in\mathbb{F}_{2}^{n}$. The expressions in parentheses on the right-hand side are all constants since $Q$ is quadratic. By the Leibniz rule (60), it suffices to show that
@@
or equivalently that $d^{4}\binom{R}{2}=\mathrm{Sym}^{2}(d^{2}Q_{1})$. But this follows from [26, Lemma 6.4, Example 6.5] (or from several direct applications of the Leibniz rule (60) using the identity
@@
for any $F\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}/4\mathbb{Z}$ and $h\in\mathbb{F}_{2}^{n}$, as well as the identities $R=Q_{1}\mod 2$, $d^{3}Q_{1}=0$, and $d^{4}R=0$). ∎
@@
Let $1\leq r\leq 5$. Then $X_{5,r}$ is an ergodic $2$-homogeneous $5$-step nilspace, and the projection map $\pi\colon X_{5,r}\to X_{2}$ given by $\pi(q,s)\coloneqq q$ for $q\in X_{2}$ and $s\in\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$ is a nilspace morphism.

### v3

@@ 1.2. Acknowledgments
OS was supported by NSF grant DMS-1926686 and ISF grant 3056/21. Over the course of this research, TT was supported by a Simons Investigator grant, the James and Carol Collins Chair, the Mathematical Analysis & Application Research Fund, and by NSF grants DMS-1764034 and DMS-2347850, and is particularly grateful to recent donors to the Research Fund. A.J. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 547294463. We thank Tim Austin, Pablo Candela, and Luka Milićević for corrections and comments. The authors are also particularly grateful to the anonymous referee for a very careful reading of the manuscript and for many helpful suggestions that improved it.
@@ 2. A characterization of coboundaries on $\mathbb{F}_{2}^{n}$
where $h_{i}=(h_{i}^{(1)},h_{i}^{(2)})\in G$, one can verify that $\rho$ is a $2$-cocycle on the elementary abelian $2$-group $G$ that does not obey (7). This $2$-cocycle will be related to a non-trivial (but now $2$-homogeneous) 55-cocycle on the degree $2$ filtered abelian group ${\mathcal{D}}^{2}(\mathbb{F}_{2}^{2})$ that we will construct in the next section.
@@
A notable special case of this theorem occurs when $\rho_{h_{1},\dots,h_{k}}$ is constant for each $h_{1},\dots,h_{k}$, then the $2$-homogeneous $k$-cocycle $\rho$ is what is referred to as a non-classical symmetric multilinear form in [28] and a strongly symmetric multilinear form in [22], and the potential $F$ produced by this theorem is then a (non-classical) polynomial of degree $k$. This special case of Theorem 2.3 was previously established in [28, Proposition 3.5].
@@
Now suppose inductively that $k>0$ and the claim has already been proven for k1k-1. For each $h_{1}\in G$, the function $\rho_{h_{1}}\colon G^{[k]}\to\mathbb{T}$ defined by $(\rho_{h_{1}})_{h_{2},\dots,h_{k+1}}(x)\coloneqq\rho_{h_{1},\dots,h_{k+1}}(x)$ can be easily verified to be a $2$-homogeneous (k1)(k-1)-cocycle. Hence by the induction hypothesis, there exists $F_{h_{1}}\colon G\to\mathbb{T}$ such that
@@
It remains to establish (12). We prove this by a further induction on the dimension $n$. The case $n=0$ is trivial, so suppose $n\geq 1$ and that the claim has already been proven for n1n-1. Now split $G=\mathbb{F}_{2}^{n-1}\times\mathbb{F}_{2}$ and let $e=(0,1)$ be the generator for the $\mathbb{F}_{2}$ factor. The operator $\partial_{e}$ is annihilated by $1+T^{e}$ since $(1+T^{e})\partial_{e}=\partial_{2e}=0$. Also, for $k>2$, the operator $1+T^{e}=2+\partial_{e}$ maps $\mathrm{Poly}^{k-2}(G)$ to $\mathrm{Poly}^{k-3}(G)$ thanks to (64), hence from (10)
@@
Since $F_{e}-F^{\prime}_{e}$ is annihilated by $1+T^{e}$, it sums to zero on each of the $2$-element cosets of $\langle e\rangle=\{0,e\}$, and we may therefore write
@@
for some $\phi\colon G\to\mathbb{T}$. For instance, if we arbitrarily select a complementary (index two) subspace $H$ to $\langle e\rangle$ in $G$, we can set $\phi(x)\coloneqq 0$ and $\phi(x+e)\coloneqq F_{e}(x)-F^{\prime}_{e}(x)$ for $x\in H$; other choices for $\phi$ are also possible.
If we then write
@@
is a polynomial of degree k2k-2 on $G$, while
@@
for all $h\in\mathbb{F}_{2}^{n-1}\times\{0\}$. On the other hand, from (9) (now applied to $F^{\prime\prime}$) and the vanishing of $F^{\prime\prime}_{e}$, we have
$F^{\prime\prime}_{h+e}-F^{\prime\prime}_{h}\in\mathrm{Poly}^{k-1}(G)$
Thus from the $e$-invariance of $\phi^{\prime\prime}$, we see that
@@
First suppose that $\rho$ is a $k$-coboundary in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, thus $\rho=d^{k+1}F$ for some $F\colon G\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$. Then we can write $\rho=d^{k}\psi$ with $\psi\coloneqq dF$; since $2F=0$, we have $2\psi=0$, and so $\rho$ is certainly strongly $2$-homogeneous.
@@ 3. A non-trivial cocycle
There exists a strongly $2$-homogeneous 55-cocycle $\rho\colon C^{6}(X_{2})\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ on $X_{2}$ taking values in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, which is not a $5$-coboundary (when viewed as a cocycle in $\mathbb{T}$).
@@
It will be convenient to adopt the following notation from [27, Definitions 6.1, 6.3].
@@
$\quad\coloneqq\sum_{\{\{i_{1},\dots,i_{k}\},\{j_{1},\dots,j_{k}\}\}:\{1,\dots,2k\}=\{i_{1},\dots,i_{k}\}\cup\{j_{1},\dots,j_{k}\}}S(h_{i_{1}},\dots,h_{i_{k}})S(h_{j_{1}},\dots,h_{j_{k}}).$
@@
and $B*B=2\mathrm{Sym}^{2}(B)$; in particular, in characteristic two we have BB=0B*B=0. By identifying $\mathbb{F}_{2}$ with a subgroup of $\mathbb{T}$, the trilinear form in Remark 2.2 can be written
(17) $\rho_{h_{1},h_{2},h_{3}}(x)=\frac{P(\mathrm{Sym}^{2}(L_{1})*L_{2})(h_{1},h_{2},h_{3})}{2}\mod 1,$
@@
A $6$-cube in $X_{2}={\mathcal{D}}^{2}(\mathbb{F}_{2}^{2})$ can be viewed as a pair $(Q^{(1)},Q^{(2)})$, where $Q^{(1)},Q^{(2)}\colon\mathbb{F}_{2}^{6}\to\mathbb{F}_{2}$ are quadratic polynomials (cf. Definition A.18), so in particular their second derivatives can be viewed as symmetric bilinear forms $d^{2}Q^{(i)}\colon\mathbb{F}_{2}^{6}\times\mathbb{F}_{2}^{6}\to\mathbb{F}_{2}$, defined for $i=1,2$ by the formula
@@
with $e_{1},\dots,e_{6}$ the standard basis of $\mathbb{F}_{2}^{6}$; observe the analogy with the construction in (17).
@@
The function $\rho$ is clearly symmetric with respect to permutations of the indices $1,\dots,6$. If we fix the $h_{ij}$ for $1<i<j\leq 6$, then this function is linear in the remaining variables $h_{1i}$, $1<i<6$, from which it is easy to verify that $\rho$ obeys the cocycle property in Definition A.6(ii). Thus $\rho$ is a $5$-cocycle.
@@
Finally, we need to show that $\rho=d^{5}\psi$ for some $\psi\colon X_{2}^{[1]}\to\mathbb{T}$ with $2\psi$ a cubic polynomial. We let $\colon\mathbb{F}_{2}\to\mathbb{Z}/4\mathbb{Z}$ be any section of the projection map $\mod 2\colon\mathbb{Z}/4\mathbb{Z}\to\mathbb{F}_{2}$; in particular one has $0]^{2}=0\mod 4$ and $1]^{2}=1\mod 4$ regardless of the choice of section. An element of $C^{1}(X_{2})$ takes the form $(x,x+h)$ with $x,h\in\mathbb{F}_{2}^{2}$. We write $x=(x^{(1)},x^{(2)})$, $h=(h^{(1)},h^{(2)})$ and define $\psi$ by the formula
(22) $\psi(x,x+h)\coloneqq\frac{[x^{(1)}]^{2}[h^{(2)}]^{2}}{4}+\frac{x^{(1)}h^{(1)}x^{(2)}}{2}\mod 1.$
We first verify that $2\psi$ is a cubic polynomial. Since $x]^{2}=x^{2}=x\mod 2$, we have
$2\psi(x,x+h)=\frac{x^{(1)}h^{(2)}}{2}\ \mod 1.$
According to Remark A.3, a $4$-cube in $C^{1}(X_{2})$ corresponds to a $5$-cube in $X_{2}$, and by Lemma A.15, a $5$-cube in $X_{2}$ can be computed according to the formula in (57), and takes the form
@@
for some $x,h_{0},h_{i},h_{0i},h_{ij}\in X_{2}$ (cf. Example A.14). The function $d^{4}(2\psi)$ applied to this cube is then equal to
@@
It remains to show that $\rho=d^{5}\psi$. As before, from Remark A.3, Lemma A.15, and the formula in (57), a $5$-cube in $C^{1}(X_{2})$ takes the form
@@
for some $x,h_{0},h_{i},h_{0i},h_{ij}\in X_{2}$. By the definition (22) of $\psi$, the function $d^{5}\psi$ applied to this cube is the sum of
@@
for $a=1,2$. We first consider (24). The numerator $X^{(1)}(\omega)H^{(1)}(\omega)X^{(2)}(\omega)$ is quintic in the variables $\omega_{i}$ (when viewed as a function from $\mathbb{F}_{2}^{5}$ to $\mathbb{F}_{2}$ for a fixed choice of the $x,h_{0},h_{i},h_{0i},h_{ij}$). We have the identity
$\sum_{\omega\in\{0,1\}^{5}}(-1)^{5-|\omega|}\prod_{i\in T}\omega_{i}=\begin{cases}1,&\text{if }T=\{1,2,\dots,5\},\\[6.0pt] 0,&\text{otherwise.}\end{cases}$
so only monomials that use all five distinct variables $\omega_{1}\dots\omega_{5}$ survive, and the sum of the coefficients of these monomials after expanding out all the definitions can be expressed as
@@
where the sum $\sum^{*}$ ranges over the $30$ pairs of sets $\{\{a,b\},\{c,d\}\},\{e,f\}$ with $\{0,1,2,3,4,5\}=\{a,b\}\cup\{c,d\}\cup\{e,f\}$ such that 0 lies in one of $\{a,b\}$ or $\{c,d\}$.
Now consider (23). Using the easily verified identities $a+b]^{2}=[a]^{2}+[b]^{2}+2[ab]$ and $a\omega]^{2}=[a]^{2}\omega$ for $a,b\in\mathbb{F}_{2}$ and $\omega\in\{0,1\}$ (and noting that the map $a\mapsto 2[a]$ is an additive homomorphism), we can expand out
$H^{(2)}(\omega)]^{2}=\left[h^{(2)}_{0}+\sum_{i=1}^{5}h^{(2)}_{0i}\omega_{i}\right]^{2}=[h^{(2)}_{0}]^{2}+\sum_{i=1}^{5}[h^{(2)}_{0i}]^{2}\omega_{i}+2[Q(\omega)]$
@@
$X^{(1)}(\omega)]^{2}=[x^{(1)}]^{2}+\sum_{i=1}^{5}[h^{(1)}_{i}]^{2}\omega_{i}+\sum_{1\leq i<j\leq 5}[h^{(1)}_{ij}]^{2}\omega_{i}\omega_{j}+2[R(\omega)]$
@@
The product $X^{(1)}(\omega)]^{2}[H^{(2)}(\omega)]^{2}$ is then quintic (the product of $2Q$ and $2R$ would be sextic, but vanishes modulo $4$), and the alternating sum $\sum_{\omega\in\{0,1\}^{5}}(-1)^{5-|\omega|}$ is then extracting the $\omega_{1}\dots\omega_{5}$ coefficient, which can only arise from the terms
@@
in the numerator. Inspecting the quartic terms of $R(\omega)$, we conclude that (23) is of the form
@@
where the sum $\sum^{**}$ ranges over the $15$ pairs of sets $\{\{a,b\},\{c,d\}\},\{e,f\}$ with $\{0,1,2,3,4,5\}=\{a,b\}\cup\{c,d\}\cup\{e,f\}$ such that 0 does not lie in either $\{a,b\}$ or $\{c,d\}$. Summing (25), (26), we obtain the claim. This concludes the proof of Theorem 3.1.
@@ 4. Two key nilspaces
We now use the non-trivial cocycle $\rho$ introduced in the previous section to construct a family of finite $5$-step nilspaces $X_{5,r}$ for $1\leq r\leq 5$ that will play a key role in our counterexamples. To prove our main result in Theorem 1.6 we will use the larger and more complicated nilspace $X_{5,5}$, however in Proposition 4.5 below we obtain a weaker counterexample with significantly less effort using the smaller and simpler nilspace $X_{5,1}$.
@@
We first show existence. Since $\rho$ is a strongly $2$-homogeneous 55-cocycle, it is not difficult to see that the pullback $Q^{*}\rho$ is also. Hence by Theorem 2.6, $Q^{*}\rho$ is a $5$-coboundary in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, thus there exists $S\colon\mathbb{F}_{2}^{n}\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ such that $Q^{*}\rho=d^{6}S$, which is precisely the condition (1). Since $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ is contained in $\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$, we have obtained an $n$-cube $(Q,S)$ in $X_{5,r}$ as required.
@@
In fact, with the specific choice of cocycle we have constructed, we can explicitly describe the coset in Lemma 4.1.
@@
Let the notation and hypotheses be as in Lemma 4.1. Write $Q=(Q_{1},Q_{2})$, thus $Q_{1},Q_{2}\colon\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ are classical quadratic polynomials. Let $R\in\mathrm{Poly}^{3}(\mathbb{F}_{2}^{n}\to\mathbb{Z}/4\mathbb{Z})$ be a cubic polynomial such that $2\frac{R}{4}=\frac{Q_{1}}{2}\mod 1$ (or equivalently that $R=Q_{1}\mod 2$); the existence of such a polynomial follows from (64). Then the coset of $S$ in Lemma 4.1 is equal to
@@
where (as in Section 1.3) $\binom{a}{2}\in\mathbb{F}_{2}$ is equal to $1$ when $a=2,3\mod 4$ and 0 for $a=0,1\mod 4$.
@@
for $x\in\mathbb{F}_{2}^{n}$ and $\vec{h}=(h_{1},\dots,h_{6})\in(\mathbb{F}_{2}^{n})^{6}$. By (20), it suffices to show that
(28) $\partial_{h_{1}}\dots\partial_{h_{6}}\left(\binom{R}{2}Q_{2}\right)=\sum_{\{\{a,b\},\{c,d\}\},\{e,f\}:\{1,\dots,6\}=\{a,b\}\cup\{c,d\}\cup\{e,f\}}(\partial_{h_{a}}\partial_{h_{b}}Q_{1})(\partial_{h_{c}}\partial_{h_{d}}Q_{1})(\partial_{h_{e}}\partial_{h_{f}}Q_{2})$
for all $h_{1},\dots,h_{6}\in\mathbb{F}_{2}^{n}$. The expressions in parentheses on the right-hand side are all constants since $Q$ is quadratic. Iteratively applying the Leibniz rule (63) and since $\partial_{h}\partial_{k}=\partial_{k}\partial_{h}$, we have
$\partial_{h_{1}}\cdots\partial_{h_{6}}(P_{1}P_{2})=\sum_{\begin{subarray}{c}S,T\subseteq\{1,\dots,6\}\\ S\cup T=\{1,\dots,6\}\end{subarray}}\biggl{(}\prod_{i\in S}\partial_{h_{i}}P_{1}\biggr{)}\biggl{(}\prod_{j\in T}\partial_{h_{j}}P_{2}\biggr{)}.$
Since $\binom{R}{2}$ is a quartic [^7] One can verify this using the identity $\binom{a+b}{2}\equiv\binom{a}{2}+\binom{b}{2}+\bigl{(}(a\bmod 2)\,(b\bmod 2)\bigr{)}\pmod{2}$. and $Q_{2}$ is a quadratic, it follows from the previous identity that in order to show (28) it suffices to show that
@@
or equivalently that $d^{4}\binom{R}{2}=\mathrm{Sym}^{2}(d^{2}Q_{1})$. But this latter identity was established in [27, Example 6.5] as a consequence of [27, Lemma 6.4]. (Alternatively, one can establish the identity $d^{4}\binom{R}{2}=\mathrm{Sym}^{2}(d^{2}Q_{1})$ by several direct applications of the Leibniz rule (63) using the identity
@@
for any $F\colon\mathbb{F}_{2}^{n}\to\mathbb{Z}/4\mathbb{Z}$ and $h\in\mathbb{F}_{2}^{n}$, as well as the identities $R=Q_{1}\mod 2$, $d^{3}Q_{1}=0$, and $d^{4}R=0$.) ∎
@@
Let $1\leq r\leq 5$. Then $X_{5,r}$ is an ergodic $2$-homogeneous 55-step nilspace, and the projection map $\pi\colon X_{5,r}\to X_{2}$ given by $\pi(q,s)\coloneqq q$ for $q\in X_{2}$ and $s\in\frac{1}{2^{r}}\mathbb{Z}/\mathbb{Z}$ is a nilspace morphism.